OpenStudy (anonymous):

1. Solve for x. 4(x – b) = x (1 point) A.b=4/3x B.b=3/4x C.x=4/3b D.x=3/4b

OpenStudy (shaik0124):

4x-4b=x now subtract x on both sides

OpenStudy (anonymous):

x = (4*b)/3

OpenStudy (anonymous):

4(x-b)=x 4x-4b=x 3x-4b=0 3x=4b x=4/3 b